# MINDS. JWST-MIRI Observations of a Spatially Resolved Atomic Jet and Polychromatic Molecular Wind Toward SY Cha

### v1

@@ Abstract
The removal of angular momentum from protostellar systems drives accretion onto the central star and may drive the dispersal of the protoplanetary disk. Winds and jets can contribute to removing angular momentum from the disk, though the dominant process remain unclear. To date, observational studies of resolved disk winds have mostly targeted highly inclined disks. We report the detection of extended H2 and [Ne II] emission toward the young stellar object SY Cha with the JWST Mid-InfraRed Instrument Medium Resolution Spectrometer (MIRI-MRS). This is one of the first polychromatic detections of extended H2 toward a moderately inclined, , Class II source. We measure the semi-opening angle of the H2 emission as well as build a rotation diagram to determine the H2 excitation temperature and abundance. We find a wide semi-opening angle, high temperature, and low column density for the H2 emission, all of which are characteristic of a disk wind. These observations demonstrate MIRI-MRS’s utility in expanding studies of resolved disk winds beyond edge-on sources.
@@ 1 Introduction
Outflows are variously classified as winds and jets based on the degree of collimation and outflow velocity. Winds launched by the disk can be driven either by high-energy radiation from the central star: photoevaporative winds, or by the stellar and/or disk magnetic field: magnetohydrodynamic (MHD) winds; see Pascucci et al. (2023) for a recent review. Photoevaporative winds are expected to be launched from radii greater than a tenth of the gravitational radius but small enough for the disk to intercept a large amount of high energy radiation from the central star, and to be relatively slow, with velocities 10 km s-1 (Hollenbach et al., 1994; Liffman, 2003; Clarke & Alexander, 2016). In contrast, MHD winds can be launched from a large range of radii, and have larger initial velocities, and thus larger vertical extents (Blandford & Payne, 1982; Lesur, 2021). Which of these processes is the dominant form of mass removal remains unclear (Hartmann et al., 2016).
Forbidden line emission from atomic ions can trace both winds and jets (Pascucci et al., 2007; van Boekel et al., 2009; Pascucci et al., 2020; Flores-Rivera et al., 2023; Tychoniec et al., 2024; Bajaj et al., 2024; Federman et al., 2024) Extended emission from ro-vibrational as well as pure rotation lines of H2 has been observed toward a number of young stellar objects (YSOs) and is often attributed to either an MHD disk wind or shocks in the outflow cavity (Beck et al., 2008; Melnikov et al., 2023; Arulanantham et al., 2024; Delabrosse et al., 2024; Tychoniec et al., 2024).
@@ 3 Analysis
The [Ne II] emission originates primarily from the region around the central source. Additionally, extended emission is clearly seen on either side of the central source, roughly perpendicular to the protoplanetary disk (Orihara et al., 2023; Ginski et al., 2024). This extended emission is strongest in the spectral images adjacent to the central [Ne II] frame at 12.814 m(see Appendix LABEL:app:channel). The blue- and red-shifted emission is binned with velocity offset of 117 km s-1, extending roughly 3″ from the central source in opposite directions. The blue-shifted emission appears brighter, as expected for a bipolar outflow where the light from the red-shifted lobe is subject to more extinction along the line of sight.
@@ 3.1 Molecular Hydrogen Rotation Diagram
The H2 emission could potentially originate from a wind or from the hot surface of the disk. In either case, the range of expected temperatures and H2 densities for our observed lines are expected to be in LTE (Le Bourlot et al., 1999). For a population of molecules in LTE, the number of molecules at each upper state excitation level, , is given by
@@
| Flux Density | ||||||
| Transition | Line Center | [erg s-1 cm-2 str-1] | ||||
| [µm] | [K] | (0-2″) | (0-0.5″) | (0.5-2.0″) | ||
| (1) | 17.0348 | 1015.1 | 21 | |||
| (2) | 12.2786 | 1681.6 | 9 | |||
| (3) | 9.66491 | 2503.7 | 33 | |||
| (4) | 8.02504 | 3474.5 | 13 | |||
| (5) | 6.90951 | 4586.1 | 45 | |||
| (6) | 6.10856 | 5829.8 | 17 | |||
| (7) | 5.51118 | 7196.7 | 57 | |||
| (8) | 5.05306 | 8677.1 | 21 | |||
Note. — and values from Roueff et al. (2019)
@@ 3.2 Atomic Ions
We determine the total flux from [Ne II], [Ne III], and [Ar II] within the central aperture, which contains primarily emission from the disk, by fitting a Gaussian to each line. As the [Ne III] overlaps with a nearby water line, two Gaussians are fit, though only the Gaussian centered on the [Ne III] is used to determine the flux, see Section B. Analysis of ions within the disk can be used to constrain the high-energy component of the stellar radiation field impinging on the disk, which in turn can drive the disk wind. In particular, the ratio of [Ne II]/[Ne III] differentiates between ionization from hard EUV photons and soft EUV and/or X-rays while [Ne II]/[Ar II] differentiates between hard X-rays and soft X-rays or EUV (Hollenbach et al., 2009). However, for SY Cha the [Ne II] emission clearly includes a substantial jet component, which can be driven by locally produced UV emission from shocks (e.g., Hollenbach, 1997) and is thus not a reliable tracer of the stellar radiation field. Higher spectral resolution observations (e.g., Pascucci et al., 2020) would be needed to separate out the jet emission from any disk emission.
@@ 4.1 Extended Molecular Hydrogen Emission
Previous studies of JWST-MIRI extended H2 emission in similar systems report narrower semi-opening angles of 9∘ to 35∘ and similar to higher temperatures, attributed to a scattered MHD disk wind (Arulanantham et al., 2024; Delabrosse et al., 2024). These systems are more inclined than SY Cha and the extended H2 was observed in the (2) and (1) transitions, which do not appear extended for SY Cha. Interestingly, Delabrosse et al. (2024) find the red-shifted wind in DG Tau B has a very narrow semi-opening angle, while the semi-opening angle of the blue-shifted emission is much wider. Based on our derived temperature and semi-opening angle, we conclude the extended H2 emission in SY Cha is likely from a disk wind. This analysis of a spatially extended molecular wind toward a moderately inclined protoplantetary disk, along with several previous studies (e.g., Güdel et al., 2018; Booth et al., 2021; Delabrosse et al., 2024), demonstrates that studies of resolved disk winds need not focus on edge-on systems.
@@ 4.2 Wind Mass Loss Rate
Assuming a typical wind velocity of 10 km s-1 (Alexander et al., 2014; Tabone et al., 2020) we can convert our derived hot H2 mass into a molecular wind mass loss rate, following the method of Arulanantham et al. (2024). We assume a maximum height above the midplane of 271 au based on the extent of the (5) emission in Figure 1 and assume an inclination angle of 51.1∘. This results in a wind mass loss rate of M⊙ yr-1, an order of magnitude lower than the mass accretion rate onto the central star reported in Table 1 and far below most reported molecular wind mass loss rates (Louvet et al., 2018; de Valon et al., 2020; Tabone et al., 2020; Arulanantham et al., 2024). However, we note that SY Cha is a transition disk with a low mass accretion rate. The disk is presumably more evolved than previously characterized systems, which could explain the lower wind mass loss rate. Additionally, the molecular wind component is likely only a fraction of the total mass in the wind, with a significant fraction of the mass in neutral and ionized atomic gas (Nakatani et al., 2018; Wang et al., 2019; Komaki et al., 2021; Sellek et al., 2024). Finally, as MHD winds are driven by accretion energy while photoevaporative winds are thermally driven (Lesur et al., 2023), it is possible that the extended H2 emission is from a photoevaporative wind, though this would not explain the [Ne II] emission.
@@ 4.3 Extended [Ne II] Emission
[Ne II] emission is spatially resolved, with both a blue shifted and red shifted component. The [Ne II] emission is both more extended and more collimated than the extended H2 emission. Similar results were recently reported for the edge-on source Tau 042021, where the jet was also seen in [Ne III], [Ni II], [Fe II], [Ar II], and [S III] (Arulanantham et al., 2024). These lines do not appear to be spatially extended in our observations of SY Cha, though this may be due in part to the weaker line emission. The undetected lines such as [Ni II], [Fe II], and [S III] could also be masked by stronger water lines in the central region.
The [Ne II] emission with a small velocity offset appears largely unresolved and co-spatial with the central source, while there are also clearly two lobes of emission extending away from the disk corresponding to velocity offsets of greater than 100 km s-1. High spectral resolution observations of [Ne II] toward other sources often show either a high-velocity component tracing a jet for sources with high accretion rates, or a low-velocity component tracing a wind for low accretors (e.g., Pascucci et al., 2020). Recent JWST MIRI MRS observations of [Ne II] in edge-on disks are likewise either attributed to a disk wind (e.g., Bajaj et al., 2024) or a jet (Arulanantham et al., 2024). For SY Cha, only the [Ne II] emission with a large velocity offset appears spatially extended. As the [NeII] emission is extended and collimated, the wind origin seems to be excluded We conclude that while there may be a wind component to the [Ne II] emission, the dominant component comes from a jet. SY Cha is unusual in that it appears to be is a low accretor with a [Ne II] jet.
@@ 5 Conclusions
1.
The [Ne II] emission is highly collimated, with both blue and red-shifted lobes, indicative of a jet.
2.
@@ 6 Acknowledgements
The data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/8ref-ye65 (catalog DOI).
This paper makes use of observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programs 0103.C-0470, 105.209E, 105.20HV, 105.20JB, 106.21HJ, and 108.22EE. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2018.1.00689.S ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
@@
I.K., J.K., and T.K. acknowledge funding from H2020-MSCA-ITN-2019, grant no. 860470 (CHAMELEON). E.v.D. acknowledges support from the ERC grant 101019751 MOLDISK and the Danish National Research Foundation through the Center of Excellence “InterCat” (DNRF150). A.C.G. acknowledges support from PRIN-MUR 2022 20228JPA3A “The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 “NIR-dark Accretion Outbursts in Massive Young stellar objects (NAOMY)” and Large Grant INAF 2022 “YSOs Outflows, Disks and Accretion: towards a global framework for the evolution of planet forming systems (YODA)”. T.P.R acknowledges support from ERC grant 743029 EASY. V.C. acknowledges funding from the Belgian F.R.S.-FNRS. D.G. thanks the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA). G.P. gratefully acknowledges support from the Max Planck Society.
7 Software & Facilities
JWST
@@ Appendix A Image De-convolution
Appendix B Fits to Atomic Lines
@@
| Transition | Line Center | Integrated Flux | rms | |
| (m) | 10-3(m) | 10-15(erg s-1 cm-2) | (erg s-1 cm-2 Hz-1) | |
| Ar II | 6.985 | 1.7 | 1.55 | 6.7 |
| Ne II | 12.814 | 2.6 | 5.87 | 8.5 |
| Ne III | 15.555 | 4.0 | 0.94 | 9.0 |

### v2

@@ Abstract
The removal of angular momentum from protostellar systems drives accretion onto the central star and may drive the dispersal of the protoplanetary disk. Winds and jets can contribute to removing angular momentum from the disk, though the dominant process remain unclear. To date, observational studies of resolved disk winds have mostly targeted highly inclined disks. We report the detection of extended H2 and [Ne II] emission toward the young stellar object SY Cha with the JWST Mid-InfraRed Instrument Medium Resolution Spectrometer (MIRI-MRS). This is one of the first polychromatic detections of extended H2 toward a moderately inclined, , Class II source. We measure the semi-opening angle of the H2 emission as well as build a rotation diagram to determine the H2 excitation temperature and abundance. We find a wide semi-opening angle, high temperature, and low column density for the H2 emission, all of which are characteristic of a disk wind. We derive a molecular wind mass loss rate of M⊙ yr-1, which is high compared to the previously derived stellar accretion rate of M⊙ yr-1. This suggests either that the stellar accretion and the disk wind are driven by different mechanisms or that accretion onto the star is highly variable. These observations demonstrate MIRI-MRS’s utility in expanding studies of resolved disk winds beyond edge-on sources.
@@ 1 Introduction
Outflows are variously classified as winds and jets based on the degree of collimation and outflow velocity. Winds launched by the disk can be driven either by high-energy radiation from the central star: photoevaporative winds, or by the stellar and/or disk magnetic field: magnetohydrodynamic (MHD) winds; see Pascucci et al. (2023) for a recent review. Photoevaporative winds are expected to be launched from radii greater than a tenth of the gravitational radius but small enough for the disk to intercept a large amount of high energy radiation from the central star, and to be relatively slow, with velocities 10 km s-1 (Hollenbach et al., 1994; Liffman, 2003; Clarke & Alexander, 2016). In contrast, MHD winds can be launched from a large range of radii, and have larger initial velocities, and thus larger vertical extents (Blandford & Payne, 1982; Lesur, 2021). Which of these processes is the dominant form of mass removal remains unclear (Hartmann et al., 2016). Jets are both more collimated (semi-opening angle ) and have a higher velocity ( km s-1) than winds (Frank et al., 2014). They are launched via MHD processes in the inner most regions of the disk or at the stellar surface, though the exact details are still uncertain (Tsukamoto et al., 2023).
Forbidden line emission from atomic ions can trace both winds and jets (Pascucci et al., 2007; van Boekel et al., 2009; Pascucci et al., 2020; Flores-Rivera et al., 2023; Tychoniec et al., 2024; Bajaj et al., 2024; Federman et al., 2024). Extended emission from ro-vibrational as well as pure rotation lines of H2 has been observed toward a number of young stellar objects (YSOs) and is often attributed to either an MHD disk wind or shocks in the outflow cavity (Beck et al., 2008; Melnikov et al., 2023; Arulanantham et al., 2024; Delabrosse et al., 2024; Tychoniec et al., 2024).
@@ 3 Analysis
The [Ne II] emission originates primarily from the region around the central source. Additionally, extended emission is clearly seen on either side of the central source, roughly perpendicular to the protoplanetary disk (Orihara et al., 2023; Ginski et al., 2024). This extended emission is strongest in the spectral images adjacent to the central [Ne II] frame at 12.814 m(see Appendix B). The blue- and red-shifted emission is binned with velocity offset of 117 km s-1, extending roughly 3″ from the central source in opposite directions. The blue-shifted emission appears brighter, as expected for a bipolar outflow where the light from the red-shifted lobe is subject to more extinction along the line of sight. Following the method of Pascucci et al. (2024), we determine the semi-opening angle of the blue and red-shifted [NeII] emission by fitting a Gaussian profile to cuts perpendicular to the outflow axis. This allows us to derive a FWHM. At this wavelength the pixel scale is . Any derived FWHM below this value is considered unresolved and not used to determine the semi-opening angle. We then derive the semi-opening angle from the slope of line fit to the remaining data points. The derived semi-opening angle for the blue-shifted lobe is , consistent with measurements for jets in other sources (Delabrosse et al., 2024; Arulanantham et al., 2024; Pascucci et al., 2024).
@@ 3.1 Molecular Hydrogen Rotation Diagram
The H2 emission could potentially originate from a wind or from the hot surface of the disk. For a population of molecules in LTE, the number of molecules at each upper state excitation level, , is given by
@@
| Flux Density | |||||||
| Transition | Line Center | [erg s-1 cm-2 str-1] | |||||
| [µm] | [K] | [cm-3] | (0-2″) | (0-0.5″) | (0.5-2.0″) | ||
| (1) | 17.0348 | 1015.1 | 21 | 2.4e2 | |||
| (2) | 12.2786 | 1681.6 | 9 | 1.7e3 | |||
| (3) | 9.66491 | 2503.7 | 33 | 5.5e3 | |||
| (4) | 8.02504 | 3474.5 | 13 | 2.0e4 | |||
| (5) | 6.90951 | 4586.1 | 45 | 3.9e4 | |||
| (6) | 6.10856 | 5829.8 | 17 | 1.0e5 | |||
| (7) | 5.51118 | 7196.7 | 57 | 1.5e5 | |||
| (8) | 5.05306 | 8677.1 | 21 | ||||
@@
We can use the derived column density for the hot component to check if our assumption of LTE is valid. To convert from an H2 column density to an average number density we divide the column density by the observed extent of the emission on the sky of 1.78″. This results in an average H2 number density of cm-3. In this high temperature, low density regime is expected to be in form of neutral H or H+, with n(H) more than a factor of 100 greater than n(H2) (Bron et al., 2016; Nakatani et al., 2018; Sellek et al., 2024). This makes atomic hydrogen the main collisional partner with a density of cm-3. According to the critical densities for collisions with H atoms at 1000 K given in Table 2, only the lowest energy lines for which we observe extended emission are expected to be in LTE (Le Bourlot et al., 1999). LTE assumes the population distribution for energy levels of a given species can be described by a single excitation temperature. Using LTE assumptions for a population in non-LTE can lead to a more than order of magnitude increase in the uncertainty of the derived column density (Mangum & Shirley, 2015). As such, the column density reported here should be considered tentative. More advanced modeling based analysis is needed to obtain more precise measurements (e.g, van der Tak et al., 2007; Brinch & Hogerheijde, 2010).
@@ 3.2 Atomic Ions
We determine the total flux from [Ne II], [Ne III], and [Ar II] within the central aperture, which contains primarily emission from the disk, by fitting a Gaussian to each line. As the [Ne III] overlaps with a nearby water line, two Gaussians are fit, though only the Gaussian centered on the [Ne III] is used to determine the flux, see Section C. Analysis of ions within the disk can be used to constrain the high-energy component of the stellar radiation field impinging on the disk, which in turn can drive the disk wind. In particular, the ratio of [Ne II]/[Ne III] differentiates between ionization from hard EUV photons and soft EUV and/or X-rays while [Ne II]/[Ar II] differentiates between hard X-rays and soft X-rays or EUV (Hollenbach et al., 2009). However, for SY Cha the [Ne II] emission clearly includes a substantial jet component, which can be driven by locally produced UV emission from shocks (e.g., Hollenbach, 1997) and is thus not a reliable tracer of the stellar radiation field. Higher spectral resolution observations (e.g., Pascucci et al., 2020) would be needed to separate out the jet emission from any disk emission.
@@ 4.1 Extended Molecular Hydrogen Emission
Previous studies of JWST-MIRI extended H2 emission in similar systems report narrower semi-opening angles of 9∘ to 35∘ and similar to higher temperatures, attributed to a scattered MHD disk wind (Arulanantham et al., 2024; Delabrosse et al., 2024). These systems are more inclined than SY Cha and the extended H2 was observed in the (2) and (1) transitions, which do not appear extended for SY Cha. Interestingly, Delabrosse et al. (2024) find the red-shifted wind in DG Tau B has a very narrow semi-opening angle, while the semi-opening angle of the blue-shifted emission is much wider. Pascucci et al. (2024) used JWST-NIRSpec to resolve H2 disk winds in the (9) transition for four edge-on disks. They find semi-opening angles ranging from to for the blue-shifted wind. The semi-opening angle for the (9) Tau 042021 is about , compared to the value of for (2) found by Arulanantham et al. (2024) using MIRI observations. Based on our derived temperature and semi-opening angle, we conclude the extended H2 emission in SY Cha is likely from a disk wind. This analysis of a spatially extended molecular wind toward a moderately inclined protoplantetary disk, along with several previous studies (e.g., Güdel et al., 2018; Booth et al., 2021; Delabrosse et al., 2024), demonstrates that studies of resolved disk winds need not focus on edge-on systems.
@@ 4.2 Wind Mass Loss Rate
Assuming a typical wind velocity of 10 km s-1 (Alexander et al., 2014; Tabone et al., 2020) we can convert our derived hot H2 mass into a molecular wind mass loss rate, following the method of Arulanantham et al. (2024). We assume a maximum height above the midplane of 271 au based on the extent of the (5) emission in Figure 1 and assume an inclination angle of 51.1∘. This results in a wind mass loss rate of M⊙ yr-1, an order of magnitude greater than the mass accretion rate onto the central star reported in Table 1 and in line with most reported molecular wind mass loss rates (Louvet et al., 2018; de Valon et al., 2020; Tabone et al., 2020; Arulanantham et al., 2024). We note that SY Cha is a transition disk with a low mass accretion rate. The disk is presumably more evolved than previously characterized systems, which could explain the lower wind mass loss rate. Additionally, the molecular wind component is likely only a fraction of the total mass in the wind, with a significant fraction of the mass in neutral and ionized atomic gas (Nakatani et al., 2018; Wang et al., 2019; Komaki et al., 2021; Sellek et al., 2024). However, our derived wind mass loss rate far exceeds the typical of 0.1-1 (Pascucci et al., 2023). As MHD winds are driven by accretion energy while photoevaporative winds are thermally driven, and thus do not transport significant angular momentum, (Lesur et al., 2023), it is possible that the extended H2 emission is from a photoevaporative wind, which does not relate to stellar accretion, though this would not explain the [Ne II] emission. Another possibility is that SY Cha’s stellar accretion rate is highly variable and was measured at a particularly low point. This would also explain the mid-IR variability between the three epochs of Spitzer observations and the JWST observations reported by Schwarz et al. (2024).
@@ 4.3 Extended [Ne II] Emission
[Ne II] emission is spatially resolved, with both a blue shifted and red shifted component. The [Ne II] emission is both more extended and more collimated than the extended H2 emission. Similar results from JWST-MIRI were recently reported for the edge-on source Tau 042021, where the jet was also seen in [Ne III], [Ni II], [Fe II], [Ar II], and [S III] (Arulanantham et al., 2024). These lines do not appear to be spatially extended in our observations of SY Cha, though this may be due in part to the weaker line emission. The undetected lines such as [Ni II], [Fe II], and [S III] could also be masked by stronger water lines in the central region.
The [Ne II] emission with a small velocity offset appears largely unresolved and co-spatial with the central source, while there are also clearly two lobes of emission extending away from the disk corresponding to velocity offsets of greater than 100 km s-1. High spectral resolution observations of [Ne II] toward other sources often show either a high-velocity component tracing a jet for sources with high accretion rates, or a low-velocity component tracing a wind for low accretors (e.g., Pascucci et al., 2020). Recent JWST MIRI MRS observations of [Ne II] in edge-on disks are likewise either attributed to a disk wind (e.g., Bajaj et al., 2024) or a jet (Arulanantham et al., 2024). For SY Cha, only the [Ne II] emission with a large velocity offset appears spatially extended. As the [NeII] emission is extended and collimated, the wind origin seems to be excluded. This nested morphology, with H2 tracing a wide angle wind and atomic forbidden line emission tracing a more collimated jet has been seen in several sources (Pascucci et al., 2024). We conclude that while there may be a spatially unresolved wind component to the [Ne II] emission close to the central star, the dominant extended component comes from a jet. SY Cha is unusual in that it appears to be is a low accretor with a [Ne II] jet.
@@ 5 Conclusions
The [Ne II] emission is highly collimated, with both blue and red-shifted lobes. This, as well as a narrow semi-opening angle of is indicative of a jet.
@@
We derive a molecular wind mass loss rate of M⊙ yr-1, in line with most literature value but unexpectedly high compared to the previously derived stellar accretion rate of M⊙ yr-1. This points to either a photoevaporative origin for the wind or variable accretion onto the central star.
6 Software & Facilities
JWST
@@ 7 Acknowledgements
The data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/61c6-ah42 (catalog doi: 10.17909/61c6-ah42).
@@
I.K., J.K., and T.K. acknowledge funding from H2020-MSCA-ITN-2019, grant no. 860470 (CHAMELEON). E.v.D. acknowledges support from the ERC grant 101019751 MOLDISK and the Danish National Research Foundation through the Center of Excellence “InterCat” (DNRF150). A.C.G. acknowledges support from PRIN-MUR 2022 20228JPA3A “The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 “NIR-dark Accretion Outbursts in Massive Young stellar objects (NAOMY)” and Large Grant INAF 2022 “YSOs Outflows, Disks and Accretion: towards a global framework for the evolution of planet forming systems (YODA)”. T.P.R acknowledges support from ERC grant 743029 EASY. V.C. acknowledgew funding from the Belgian F.R.S.-FNRS. D.G. thanks the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA). G.P. gratefully acknowledges support from the Max Planck Society.
@@ Appendix A Image De-convolution
Appendix B Channel Maps
Here we present the channel maps for lines with extended emission Figures 13. These maps have been continuum subtracted using the flux in adjacent channels. No de-convolution has been applied.
Appendix C Fits to Atomic Lines
@@
| Transition | Line Center | Integrated Flux | rms | |||
| [m] | 10-3[m] | [K] | [cm-2] | 10-15[erg s-1 cm-2] | [erg s-1 cm-2 Hz-1] | |
| Ar II | 6.985 | 1.7 | 2059.8 | 4.2e5 | 6.7 | |
| Ne II | 12.814 | 2.6 | 1122.0 | 6.3e5 | 5.87 | 8.5 |
| Ne III | 15.555 | 4.0 | 924.9 | 2.7e5 | 0.94 | 9.0 |